### v2

# VT-Former: A Transformer-based Vehicle Trajectory Prediction Approach For Intelligent Highway Transportation Systems

Abstract
Enhancing roadway safety and traffic management has become an essential focus area for a broad range of modern cyber-physical systems and intelligent transportation systems. Vehicle Trajectory Prediction is a pivotal element within numerous applications for highway and road safety. These applications encompass a wide range of use cases, spanning from traffic management and accident prevention to enhancing work-zone safety and optimizing energy conservation. The ability to implement intelligent management in this context has been greatly advanced by the developments in the field of Artificial Intelligence (AI), alongside the increasing deployment of surveillance cameras across road networks. In this paper, we introduce a novel transformer-based approach for vehicle trajectory prediction for highway safety and surveillance, denoted as VT-Former. In addition to utilizing transformers to capture long-range temporal patterns, a new Graph Attentive Tokenization (GAT) module has been proposed to capture intricate social interactions among vehicles. Combining these two core components culminates in a precise approach for vehicle trajectory prediction. Our study on three benchmark datasets with three different viewpoints demonstrates the State-of-The-Art (SoTA) performance of VT-Former in vehicle trajectory prediction and its generalizability and robustness. We also evaluate VT-Former’s efficiency on embedded boards and explore its potential for vehicle anomaly detection as a sample application, showcasing its broad applicability.
Index Terms:
Trajectory Prediction, Road Safety, Intelligent Transportation, Transformer, Cyber-physical Systems, Artificial Intelligence
I Introduction
In the realm of Cyber-Physical Systems (CPS) and its intersection with Computer Vision, the subject of Vehicle Trajectory Prediction (VTP) has emerged as a point of significance. VTP is utilized within various higher-level applications among which are traffic management and control [27, 52, 36, 39], accident and collision avoidance [17, 25, 40], workzone safety [27, 62, 64], route planning [50, 65], and vehicle anomaly detection [48, 26, 19, 24], etc.
The rapid advancements in AI and the widespread availability of surveillance cameras have opened up promising avenues and introduced novel techniques for advancing vehicle trajectory prediction methods [13, 20, 32]. VTP is defined as a process that utilizes the historical vehicle movement to forecast the future path, position, and behavior of a vehicle over a specific time horizon. While many studies focus on trajectory prediction from the perspective of autonomous driving and autonomous vehicles [57, 18, 6], some studies predominantly approach it from surveillance angles, including the bird’s eye view, high-angle, and eye-level perspectives [4, 27]. In this paper, we focus on the surveillance angles in the context of highway safety.
There are several challenges and requirements associated with VTP, among which three stand out. Firstly, there are real-time constraints for VTP. Many applications of trajectory prediction, such as traffic management or vehicle anomaly detection, require real-time predictions, which add complexity in terms of computational resources and speed. Balancing the trade-off between model accuracy and computational efficiency is a challenge [43, 33, 4]. Secondly, vehicles interact with each other and with pedestrians, cyclists, and other road users. Predicting how these interactions will unfold is complex, especially when combined with the uncertain dynamics of these moving objects. Another substantial challenge lies in the limited flexibility of most proposed methods, as they often lack the adaptability to generalize effectively across diverse setups primarily tailored for specific dataset conditions. Consequently, the inherent variability of environmental factors, such as camera angles and lighting conditions, can significantly impair their performance and overall robustness.
In this paper, we present VT-Former, an innovative approach that leverages the power of transformers for vehicle trajectory prediction in roadways. VT-Former also introduces Graph Attentive Tokenization (GAT) to capture the underlying social interactions among moving vehicles within the scene. It allows us to take full advantage of spatial and temporal features, resulting in a holistic and accurate approach to modeling the complex dynamics of vehicle movements in diverse and dynamic environments. Integrating these key components, GAT, and the decoder-only transformer, within VT-Former, offers a State-of-the-Art (AotA) algorithmic approach for vehicle trajectory prediction.
To demonstrate the benefits of VT-Former, we perform extensive experimentation using three CPS-relevant benchmark datasets: NGSIM [11, 12], CHD High-angle and CHD Eye-level [28] datasets, to comprehensively evaluate the performance of the VT-Former model from a multidimensional perspective. The evaluation encompassed three distinct angle variations corresponding to each dataset. VT-Former demonstrated exceptional performance, surpassing the efficacy of prior models. Notably, on the NGSIM dataset, our design achieved an Average Displacement Error (ADE) of 0.82 meters and a Final Displacement Error (FDE) of 1.86 meters, establishing a new state-of-the-art (SoTA) standard for the prediction of vehicle trajectories. Additionally, VT-Former demonstrated competitive results on both the CHD Eye-level and High-angle datasets.
To illustrate the adaptability of VT-Former for CPS applications, we assessed its performance on an embedded system. We delved into a thorough analysis, comparing the effects of using shortened historical trajectory data against extended historical horizons. This analysis revealed the model’s precision in forecasting future trajectories. Such an attribute is essential across numerous applications, especially in contexts like vehicle anomaly detection where prompt responses are paramount. Emphasizing VT-Former’s value in a vital CPS application aimed at boosting highway safety, we introduce an intuitive yet potent driving anomaly detection method. Drawing inspiration from predictive-based anomaly detection techniques in domains such as cyber security and public safety, we utilized VT-Former’s state-of-the-art accuracy to gauge its proficiency in predictive-based driving anomaly detection. We validated our approach using a dataset of anomalous driving behaviors, generated through an adversarial anomalous trajectory creation algorithm from NGSIM, highlighting its promising capability to tackle real-world issues.
Overall, the contributions of this paper can be summarized as follows:
  • Introducing VT-Former. A novel approach for vehicle trajectory prediction for highway safety applications.
  • Comprehensive assessments on benchmark datasets encompassing diverse surveillance camera angles, including bird’s-eye view, high-angle view, and eye-level view.
  • Exploring the influence of reduced observation horizons on model accuracy, underscoring its effectiveness in scenarios demanding rapid trajectory prediction.
  • Assessing the real-time efficiency of the model by executing it on an embedded platform.
  • Demonstrating the practical utility of VT-Former by implementing a straightforward driving anomaly detection algorithm and evaluating it on adversarially generated anomaly scenarios.
II Related Works
In recent years, intelligent transportation systems (ITS) have emerged as a focal point of extensive research endeavors within the scientific community [1, 53, 16, 44, 58]. Modern edge-based Cyber-Physical Systems are increasingly integrating deep learning-based prediction mechanisms [47, 4, 38]. For many CPS applications, path prediction is crucial as it helps in foreseeing future actions and locations beneficial for many downstream tasks. These predictive tools operate in real-time at local edge nodes [21, 27]. Such methodologies can be utilized for forecasting traffic densities [22, 9], improving the safety of work zones [54], and promoting autonomous vehicle technologies [55]. These predictive models also play a critical role in enabling advancements in smart monitoring and surveillance for enhanced security [49], improving pedestrian and worker safety in urban environments [42, 61], facilitating the identification and forecasting of anomalies within complex systems [41], and optimizing the design of transportation routes for increased efficiency and resource allocation [7].
Traditional VTP methods, such as those based on Long Short-Term Memory (LSTM) networks, have been explored by Alahi et al. [2]. While effective, these methods often lack the flexibility and scalability of newer architectures. Graph Neural Networks (GNNs) have emerged as a powerful tool for trajectory prediction [35, 15], capitalizing on their ability to model relational data and complex interactions between entities. Specifically, Graph Convolutional Networks (GCN) [10, 55] allows for localized feature learning on graphs, making them apt for capturing spatial relationships between vehicles. The Graph Isomorphism Network (GIN) [42] further extends this capability by distinguishing different graph structures, enabling more precise trajectory predictions by considering nuanced vehicular interactions [66, 60, 31].
In recent studies, there has been a notable shift from Long Short-Term Memory (LSTM) and Convolutional Neural Networks (CNNs) to Transformers, as introduced by Vaswani et al. [59]. Transformers employ self-attention mechanisms for the analysis of sequential data making them suitable for many applications such as computer vision [30], natural language processing [56], molecular chemistry [51], etc. Simultaneously, there have been endeavors to harness the power of Transformers in predicting trajectories within the realm of autonomous vehicles, enabling a better understanding of the temporal evolution of object movements [67, 23], yet these methods have yet to be applied in the context of highway surveillance and safety applications.
Despite the multitude of trajectory prediction algorithms proposed over the past decade, a limited number are explicitly designed for highway environments, with even fewer tailored for resource-constrained edge devices [5, 27]. Furthermore, the field of highway trajectory prediction is hampered by the scarcity of comprehensive datasets. Some recent works such as [28] have attempted to mitigate this and fill the gap. However, there remains a distinct requirement for specialized datasets tailored to highway-based applications. In general, there is a noticeable disparity between research endeavors and real-world applications in the realm of trajectory prediction based on already installed surveillance cameras. This underscores the potential for research and further development in this domain.
III VT-Former Formulation and Methodology
Figure 1: Graph Attentive Tokenization (GAT). GAT gets the past trajectory ($C_{i}$) of each vehicle ($V_{i}$) over the observation horizon ($T_{OH}$) for each vehicle and calculates the relative movement ($\Delta C_{i}$). These two components are concatenated and passed through a fully connected layer, expanding the feature maps by a factor of 2. Subsequently, we employ a Graph Network to capture intricate interactions among vehicles within the scene. The output of this Graph Network is further concatenated with the relative movement, accentuating the temporal evolution within the sequence. Finally, to infuse temporal order into the token sequences, we leverage temporal encoding.
At its core, VT-Former is a Transformer based approach for vehicle pass prediction in roadways. Transformers are renowned for their exceptional generalization capabilities [59], enabling them to discern intricate long-term patterns within data and excel in tasks featuring extensive and diverse datasets. While the original Transformer model was computationally demanding and exhibited slower performance as it had both encoder and decoder components, numerous optimizations and model adaptations have been developed to enhance real-time processing [29, 8] by choosing only encoder or decoder as the core of their models. This has facilitated the utilization of transformers in applications necessitating low-latency responses. These attributes motivated our choice of the decoder-only transformer [8] as the core component of our algorithm.
VT-Former model utilizes a temporal window of the movement of a vehicle as input. This window is precisely determined by the spatial coordinates denoting the center of the vehicle’s bounding box and extends across the entirety of the observation horizon, creating a temporal window of the movements of the vehicle along the time axis:
$C_{i}=[(x_{i},y_{i})^{t-T_{OH}+1},(x_{i},y_{i})^{t-T_{OH}+2},...,(x_{i},y_{i})^{t}]$ (1)
where $(x_{i},y_{i})^{t}$ represents the center of the $i^{th}$ vehicle’s bounding box and $T_{OH}$ is the observation horizon. The model predicts the future location of vehicles through the prediction horizon ($T_{PH}$):
$PC_{i}=[(x^{\prime}_{i},y^{\prime}_{i})^{t+1},(x^{\prime}_{i},y^{\prime}_{i})^{t+2},...,(x^{\prime}_{i},y^{\prime}_{i})^{t+T_{PH}}]$ (2)
The overview of VT-Former can be seen in Fig. 2. Our proposed design consists of two main modules; the Graph Attentive Tokenization (GAT) module and the Transformer Prediction module. In the following subsections, we will delve into the details of the proposed architecture.
Figure 2: VT-Former at a glance: our approach begins with the application of Graph Attentive Tokenization (GAT) to create enriched token sequences that encapsulate social interactions. Subsequently, these enhanced sequences are fed through the Transformer Prediction module, which consists of 8 layers of transformer decoders, each equipped with 4 heads, generating predicted trajectories. $T_{OH}$, $TS_{i}$, $T_{PH}$, and $PC_{i}$ denote the observation horizon time, token sequence of the $i^{th}$ vehicle, prediction horizon time, and the predicted trajectory, respectively.
III-A Graph Attention Tokenization
Graph Attention Tokenization (GAT) illustrated in Fig. 1 aims to extract enriched tokens that encapsulate the social interactions between the vehicles available in the scene. In order to accentuate the relative movements of the vehicles through the observation window we define the relative trajectory with respect to the first observed location as:
$\begin{split}\Delta C_{i}=[(x_{i},y_{i})^{t-T_{OH}+1}-(x_{i},y_{i})^{t-T_{OH}+1},\\ (x_{i},y_{i})^{t-T_{OH}+2}-(x_{i},y_{i})^{t-T_{OH}+1},\\ ...,(x_{i},y_{i})^{t}-(x_{i},y_{i})^{t-T_{OH}+1}]\end{split}$ (3)
$C_{i}$ and $\Delta C_{i}$ are concatenated to form the input of the fully connected layer which expands the dimension of the input features by a factor of $2$. In the next step, we leverage Graph Isomorphism Network (GIN) [66] to capture the relationships between subjects available in the scene. GINs have proven highly effective in the context of processing graph-structured data, where nodes and edges signify entities and their relationships, respectively. In our formulation, each vehicle will be considered as a node in a fully connected graph. We chose to have a fully connected graph so that the Graph Network does not have any predefined biases and all relationships are learned by training. Another benefit of having a fully connected graph is that by performing just one step graph operation, the information will be propagated over all nodes and further steps are not necessary which makes the design more efficient for CPS applications. Inspired by [4], we design separate aggregation networks for the node feature and the neighbor features. The node feature aggregator is an MLP network with one hidden layer:
$E^{\prime}_{v}=Act(W_{1}\cdot Act(W_{0}\cdot(\beta\cdot E_{v})+B_{0})+B_{1})$ (4)
where $E_{v}$ is the feature map for vehicle $v$, $\beta$ is a learnable parameter, and $W_{0}$, $B_{0}$, $W_{1}$ and $B_{1}$ are the parameters of an MLP with one hidden layer. We have used the Leaky ReLu function as an activation function.
For aggregating neighbor features, we designed a more sophisticated network equipped with an attention mechanism:
$\begin{split}E^{\prime}_{u}=Act(W_{2}\cdot E_{u}+B_{2})\\ E^{\prime\prime}_{u}=A_{s}[A_{C}(E^{\prime}_{u})\times E^{\prime}_{u}]\times[A_{C}(E^{\prime}_{u})\times E^{\prime}_{u}]+E^{\prime}_{u}\\ \text{where u}\in\mathcal{N}(v)\end{split}$ (5)
$E_{u}$ is the feature map of neighbour vehicle, $\mathcal{N}(v)$ is the set of neighbour vehicles, $W_{2}$ and $B_{2}$ are the parameters of a fully connected layer, $Act$ is Leaky ReLu function, and $A_{s}$ and $A_{c}$ are spatial attention and channel attention mechanism introduced by [63]. The attention mechanism helps the network to focus more on the features from neighbors that are most informative for predicting the future trajectory. The sum of $E^{\prime}_{v}$ and $E^{\prime\prime}_{u}$ for all neighbor nodes is again concatenated with $\Delta C_{i}$ to emphasize relative movement constructing the predictor input sequence. In order to also add temporal ordering to the tokens in the predictor input sequence, we leverage temporal encoding proposed by [59] as depicted by Fig. 1. This process is necessary due to the fact that the Transformer Predictor module in the next step is not able to grasp the order of the sequence as suggested by [59].
III-B Transformer Predictor Module
In the domain of time series data processing, Recurrent Neural Networks (RNNs) have traditionally been the primary choice. However, RNNs exhibit limitations related to their short-term memory, often constrained by challenges like vanishing gradients, which limits their ability to capture relationships between distant elements within a sequence. Several works also have tried using variations of Convolutional Neural Networks (CNNs) [4, 42] for finding temporal patterns, however, CNNs are not the preferred choice for modeling time series data because they lack the ability to effectively capture temporal dependencies, handle variable sequence lengths, and recognize long-range sequential patterns due to limited kernel size. Recently, a substantial paradigm shift has occurred with the emergence of transformers [59]. Transformers have garnered significant attention, predominantly due to their capability to capture long-range dependencies. Originally designed for Natural Language Processing (NLP) tasks, transformers have been adopted across a vast number of domains, including but not limited to computer vision, image and video processing, etc.
Hence, we have adopted a transformer-based architecture for more accurate vehicle trajectory prediction. The attention score inside the transformer layers is calculated using three learnable matrices Query (Q), Key (K), and Value (V) as described in [59]:
$\operatorname{Attention}(Q,K,V)=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (6)
where $d_{k}$ is the dimensionality of the Key matrix. Multiple attention heads are employed in transformers making them capable of attending to various input patterns. In our network, we have opted for a configuration with four attention heads. Following the attention mechanism, a Feed Forward Network (FFN) with a single hidden layer is utilized:
$\operatorname{FFN}(x)=\operatorname{Act}\left(x\cdot W_{3}+b_{3}\right)\cdot W_{4}+b_{4}$ (7)
where $W_{3}$, $b_{3}$, $W_{4}$, and $b_{4}$ are the parameters of the two fully connected layers. The Transformer Predictor Module consists of 8 decoder layers stacked on top of each other. We chose the dimension of the Feed Forward Network (FFN) to be a size of 256. Within our design, as detailed in Sec. III-A, we incorporated inter-vehicle attention in the tokenization process, although temporal attention among tokens was not considered. The multi-head attention mechanism in transformers is an ensemble of distinct attention mechanisms, with each head specializing in particular patterns. This adaptability empowers transformers to effectively capture a wide spectrum of data dependencies in the temporal dimension, ultimately improving the accuracy of predicted trajectories. We choose an autoregressive strategy where the model generates the sequence of predicted locations one step at a time in an ordered manner, where each prediction is influenced by the previous predictions leading to more accurate prediction. To prevent the Transformer Predictor from peeking into future information, we employ future masking while training. The masking procedure is particularly important to ensure the causality of the system in parallel training mode.
IV Experimental Setup
All trainings have been conducted on a server equipped with 2x EPYC 7513 processors and 3x Nvidia A6000 GPUs with 256GB of memory.
IV-A Datasets
IV-A1 NGSIM
The NGSIM dataset, specifically the US-101 [12] and I-80 [11] subsets, offer a rich source of real-world traffic scenarios and driver behaviors. As suggested by the name, this dataset is gathered from two distinct locations namely southbound US 101 in Los Angeles, CA, and eastbound I-80 in Emeryville, CA. With millions of data samples captured in a bird’s-eye view at 10 Hz, it provides a comprehensive perspective of the surrounding environment. NGSIM data is originally collected from cameras positioned on buildings near the freeways, and it is then converted into a bird’s-eye format. The bounding box and tracking information are all hand-annotated and thus the noise is minimal in this dataset. This holds particular significance in practical, real-world scenarios, especially when dealing with pre-installed surveillance camera systems where manually annotated data is absent. Consequently, transitioning from training on hand-annotated data to operational deployment in the real world may result in a degradation in the model’s performance.
IV-A2 Carolina’s Highway Dataset
Carolina’s Highway Dataset (CHD) [28] is a comprehensive collection of vehicle trajectory data extracted from videos recorded in various locations across North Carolina and South Carolina, USA. It includes two distinct viewpoints: eye-level (CHD Eye-level) and high-angle (CHD High-angle), each included in a separate dataset. CHD not only provides trajectory information but also includes annotations for multi-vehicle detection and tracking. The data was recorded in full HD 60Hz at eight different locations, each chosen to capture a wide range of traffic patterns, vehicle behaviors, and lighting conditions. CHD offers a diverse perspective on incoming traffic, making it a valuable resource for understanding real-world highway scenarios. The annotations for CHD Eye-level and CHD High-angle datasets are derived through the application of deep learning models. Owing to the machine-based annotation process, these datasets exhibit a greater amount of noisy data in comparison to the NGSIM dataset. However, in contrast to NGSIM, these machine-generated annotations readily align with conditions and data obtainable in real-world settings.
IV-B Metrics
IV-B1 Average Displacement Error (ADE)
is a metric commonly used in the field of trajectory forecasting and motion prediction. It quantifies the accuracy of predicted trajectories for moving objects, such as vehicles or pedestrians by measuring the average Euclidean distance between the ground truth coordinates ($x^{t},y^{t}$) and the predicted coordinates ($x^{\prime t},y^{\prime t}$) across a defined prediction horizon ($T_{PH}$) considering all available $N$ vehicles in the scene:
$ADE=\frac{\sum_{i=1}^{N}\sum_{t=1}^{T_{PH}}\sqrt{\left(x^{t}-x^{\prime t}\right)^{2}+\left(y^{t}-y^{\prime t}\right)^{2}}}{N\times T_{PH}}$ (8)
In essence, it gauges how closely the predicted trajectory aligns with the actual trajectory, providing a numerical assessment of the model’s forecasting performance. A lower ADE value indicates more accurate predictions, while a higher value suggests less accurate or less precise trajectory forecasts. ADE is a fundamental measure in evaluating the quality and reliability of trajectory prediction models.
IV-B2 Final Displacement Error (FDE)
quantifies the Euclidean distance between the ground truth final coordinates ($x^{T_{PH}},y^{T_{PH}}$) and the predicted final coordinates ($x^{\prime T_{PH}},y^{\prime T_{PH}}$) averaged over all available $N$ vehicles:
$FDE=\frac{\sum_{i=1}^{N}\sqrt{\left(x^{T_{PH}}-x^{\prime T_{PH}}\right)^{2}+\left(y^{T_{PH}}-y^{\prime T_{PH}}\right)^{2}}}{N}$ (9)
In essence, FDE assesses how accurately the model can forecast the precise destination or concluding point of an object’s trajectory. A smaller FDE value indicates more accurate predictions, signifying that the model successfully anticipates the final position, while a larger FDE value indicates less precise forecasting, implying a greater deviation from the actual endpoint.
IV-B3 Root Mean Squared Error (RMSE)
measures the square root of the average of the squared differences between the ground truth trajectory and the predicted trajectory at time step $t$ over all available $N$ vehicles:
$RMSE^{t}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left[\left(x^{t}-x^{\prime t}\right)^{2}+\left(y^{t}-y^{\prime t}\right)^{2}\right]}$ (10)
A smaller RMSE indicates that the model’s predictions are closer to the actual values, while a larger RMSE signifies a greater level of error in the predictions.
IV-C Training Strategy and Hyperparameters
In accordance with the established conventions of prior research, our study maintains an observation horizon spanning 15 timesteps (equivalent to 3 seconds) and a prediction horizon extending to 25 timesteps (equivalent to 5 seconds) for all the datasets under consideration. Given the disparate recording rates of these datasets, a standardization process is undertaken to achieve comparability with previous research, necessitating the down-sampling of both datasets to a common frame rate of 5 frames per second.
Furthermore, aligned with established precedents, any instances where the complete trajectory data of a vehicle within a specified time window is unavailable are excluded from the datasets. This practice serves as a filtering and denoising mechanism. It is essential to acknowledge that, as discussed in Sec. IV-A, due to the relatively noisier nature of annotations within the CHD High-level and CHD Eye-level datasets in comparison to the NGSIM dataset, a notably higher volume of data points are omitted in the former datasets as part of this data refinement process.
Throughout all training sessions, the models undergo 80 training epochs, with a learning rate of 0.01, a weight decay of 0.0005, and a dropout rate of 0.2 consistently applied using Adam optimizer with batch size set to 16.
V Results
TABLE I: Comprasion with SotA methods on NGSIM [12, 11]. The results are in meters.
RMSE
Model ADE FDE 1s 2s 3s 4s 5s
CS-LSTM [14] 2.29 3.34 0.61 1.27 2.09 3.10 4.37
GRIP++ [34] 2.01 3.25 0.38 0.89 1.45 2.14 2.94
STA-LSTM [37] 1.89 3.16 0.37 0.98 1.71 2.63 3.78
DeepTrack [27] 2.01 3.21 0.47 1.08 1.83 2.75 3.89
Pisghu [4] 0.88 1.96 0.15 0.46 0.82 1.25 1.74
VT-Former 0.82 1.86 0.13 0.42 0.77 1.17 1.64
TABLE II: Comparison with SotA methods on CHD [28] High-angle dataset. The results are in pixels.
RMSE
Model ADE FDE 1s 2s 3s 4s 5s
S-STGCNN [45] 31.87 98.46 9.74 21.83 29.01 42.34 82.14
GRIP++ [34] 36.32 100.89 3.40 6.67 14.32 28.02 123.04
Pisghu [4] 18.33 61.92 4.04 7.48 13.99 24.30 51.51
VT-Former 25.95 87.21 7.60 17.35 22.90 27.97 66.39
TABLE III: Comparison with SotA methods on CHD [28] Eye-level dataset. The results are in pixels.
RMSE
Model ADE FDE 1s 2s 3s 4s 5s
S-STGCNN [45] 24.33 95.22 4.32 9.15 15.93 29.05 68.32
GRIP++ [34] 44.27 129.58 4.42 12.86 24.31 35.04 145.17
Pisghu [4] 37.99 123.69 4.98 13.58 26.61 50.31 106.45
VT-Former 34.88 100.59 6.71 17.24 28.70 45.86 82.00
V-A Comparison with State-of-the-Art Methods
In this section, we conduct a comparative analysis between VT-Former and other trajectory prediction methods, leveraging three extensive highway datasets discussed in Sec. IV-A: NGSIM [12, 11], CHD High-angle [28] and CHD Eye-level.
The results, presented in Tab. I, highlight performance metrics ADE, FDE, and RMSE on the NGSIM dataset. Values reported for GRIP++ [34] and STA-LSTM [37] are reported from [27]. VT-Former sets a new SotA on the NGSIM dataset, showcasing improvements of $6.8\%$ and $5.1\%$ in terms of ADE and FDE, respectively, surpassing the previous best-performing methods. In terms of RMSE, our model consistently demonstrates reduced errors, affirming its precision in trajectory prediction and establishing them as the leading choice for accurate trajectory prediction. The NGSIM dataset, comprising over 100,000 frames of data, offers a substantial volume of information, allowing the multi-head attention mechanism in our Transformer Predictor Module to be effectively trained. This, in turn, facilitates the model in capturing and learning intricate temporal patterns. Both CHD datasets are recently released, and as such, the majority of SotA models have not been evaluated on this specific dataset. To bridge this gap, we have used the available code repositories of leading SotA models, reconfiguring them for training and evaluation on the CHD datasets. As illustrated in Tab. II and Tab. III, our model ranks as the second-best performer in both the CHD High-angle and Eye-level datasets. This highlights its remarkable adaptability across a spectrum of environmental conditions and viewing angles, setting it apart from other models that tend to experience large performance degradation when transitioning between different settings. These outcomes underscore the synergistic effect of Graph Attentive Tokenization and the Transformer Predictor, which, when combined, capture complementary features, resulting in a versatile model suitable for deployment in various settings.
V-B Observation Horizon Analysis
TABLE IV: VT-Former results using different observation horizons on NGSIM [12, 11] dataset. The results are in meters.
RMSE
$\mathbf{T_{OH}}$ ADE FDE 1s 2s 3s 4s 5s
1s 0.82 1.87 0.13 0.41 0.76 1.17 1.65
2s 0.88 1.97 0.15 0.45 0.82 1.26 1.75
3s 0.82 1.86 0.13 0.42 0.77 1.17 1.64
TABLE V: VT-Former results using different observation horizons on CHD High-angle dataset. The results are in pixels.
RMSE
$\mathbf{T_{OH}}$ ADE FDE 1s 2s 3s 4s 5s
1s 25.33 88.99 5.67 12.96 19.12 29.83 70.72
2s 25.90 87.90 6.40 14.62 20.56 29.44 70.05
3s 25.95 87.21 7.60 17.35 22.90 27.97 66.39
TABLE VI: VT-Former results using different observation horizons on CHD [28] Eye-level dataset. The results are in pixels.
RMSE
$\mathbf{T_{OH}}$ ADE FDE 1s 2s 3s 4s 5s
1s 21.86 66.28 5.42 12.69 18.81 26.67 53.05
2s 27.44 85.45 5.19 12.90 21.38 35.09 68.60
3s 34.88 100.59 6.71 17.24 28.70 45.86 82.00
In this section, we conduct a comprehensive investigation into the influence of different observation horizons to discern the optimal configuration for each dataset. The observation horizon is varied, ranging from 1 second to 3 seconds. Subsequently, we train and assess the performance of VT-Former on the NGSIM [12, 11], CHD High-angle [28], and CHD Eye-level [28] datasets. The results are presented in Tab. IV, Tab. III and Tab. V for the respective datasets. This specific assessment is especially important for applications that require faster prediction, such as collision avoidance or anomaly detection.
In the context of the NGSIM dataset, we have observed relatively minor variations in the accuracy of predicted trajectories when employing different observation horizons. These variations are at most $7.3\%$ and $5.3\%$ degradation in terms of ADE and FDE, respectively, as compared to the conventional 3-second observation setup. This particularly shows that even with a lower observation horizon, VT-Former is still able to accurately predict future trajectories. In the CHD High-angle dataset, a similar trend is observed, with the accuracy of predicted trajectories displaying consistency and minimal disparities among various observation horizons.
However, the results on the CHD Eye-level dataset, as depicted in Tab. VI, offer distinct insights. Here, it is evident that a shorter observation horizon is significantly more advantageous. Specifically, when deploying a 1-second observation horizon, we observe a remarkable enhancement of $37.3\%$ in FDE and $34.1\%$ in ADE compared to 3-second results. This outcome underscores the practicality of employing shorter observation horizons in the eye-level setup. This phenomenon can be attributed to the characteristics of the eye-level angle, wherein older observed frames are positioned farther from the camera. Consequently, a higher percentage of occlusions is encountered. Given that the annotations in the CHD datasets are extracted using deep learning algorithms, the increased occlusion contributes to a degradation in the quality of the extracted annotations, resulting in less accurate input data for trajectory prediction. Thus, a lower observation horizon translates to lower noise, and finally a more accurate prediction.
In summary, the empirical evidence presented in this study shows that, in practical, real-world high-angle and eye-level point-of-view scenarios marked by elevated occlusion rates among the detected vehicles, the adoption of shorter observation horizons is an effective strategy for decreasing the negative impact of noisy data. This approach not only leads to more accurate trajectory prediction outcomes but also offers the advantage of expeditious prediction. By reducing the temporal extent of observation, the model can effectively filter out the adverse effects of occlusion, which normally is seen in vehicles that are further from the camera, thus enhancing the precision of its predictions and enabling it to respond more swiftly. This finding underscores the practical utility of optimizing observation horizons in trajectory prediction models, particularly in applications where the timely and precise forecasting of object movements is important.
V-C Real-time Evaluation
Evaluating the real-time capabilities of VTP algorithms on embedded platforms is essential in ensuring the safety and reliability of these systems, as in many cases such as vehicle anomaly detection, they must make a split-second decision. By running trajectory prediction algorithms on an embedded board, developers can assess their performance in resource-constrained environments, and meet the low-latency requirements necessary to react promptly to changing conditions on the road. This process is not only important from the safety and efficiency perspectives but also helps bring cost-effective and energy-efficient technologies to a broader range of consumers.
(a) ROC curve with an AUC of 0.71
(b) PR curve with an AUC of 0.82
(c) FNR/FPR plots with EER of 0.32
Figure 3: Anomaly detection performance using VT-Former with 3-second input and 5-second prediction window.
VT-Former has been tested on Nvidia Jetson Orin AGX. The results can be seen in Tab. VII. As expected and discussed in Sec. IV-C it can be seen that the vehicles per frame of the NGSIM [11, 12] is higher than CHD Eye-level and CHD High-angle and the latency of NGSIM is significantly lower than each of the CHD datasets. As shown by the results, the latency of the model is 5.24, 69.89, and 48.09 milliseconds for NGSIM, CHD Eye-level, and CHD High-angle, respectively, indicating that the process can be done in real time. The results showcase that VT-Former can be adopted in the settings that require real-time processing, and where the speed of prediction and processing is of importance.
TABLE VII: VT-Former latency results on Jetson Orin AGX. The results are gathered using the test set of each dataset.
Dataset
Point
of View
Vehicles
per Frame
Latency
NGSIM [11, 12] Bird’s-eye 191 5.24
CHD [28] Eye-level 14 69.89
CHD [28] High-angle 21 48.09
VI VT-Former Usecase: Driving Anomaly Detection
Driving anomaly detection in an unsupervised setting refers to the process of identifying unusual or abnormal behavior or patterns in the operation of vehicles without the need for labeled training data. As anomaly detection is an open-set problem, with unsupervised setting, the model learns to predict the normal behavior in the training, and later in test and validation, when there is an anomalous behavior, the model’s prediction deviates from what is shown in the ground truth [3]. This type of anomaly detection is typically used for various purposes, such as ensuring the safety and security of vehicles, preventing accidents, and optimizing vehicle performance. The application of prediction-based unsupervised anomaly detection methodologies is a common practice in the field of pedestrian anomaly detection, particularly when focusing on pose-based anomaly detection [41, 3, 46]. In the context of driving anomaly detection, we extend these fundamental principles, as the trajectory of vehicles exhibits a close correlation with the pose exhibited by pedestrians.
In trajectory prediction, precision is imperative for constructing robust world models that anticipate naturalistic behaviors. Currently, VT-Former stands as the SotA on the NGSIM dataset [11, 12], exemplifying an accurate representation of these behaviors. The VT-Former model, trained on the NGSIM dataset, is capable of predicting normal driving behaviors. However, since it hasn’t been exposed to abnormal data during training it might struggle to accurately anticipate abnormal trajectories, resulting in significant deviations from ground truth. Hence, a higher ADE can serve as an effective metric for detecting anomaly scores.
While highways and roads are considered to be risky environments, there is a lack of proper anomalous driving datasets in this area. One of the main and valid reasons is that anomalies such as accidents are too dangerous for artificial creation. As such, normally an adversarial anomaly generator is utilized for assessment. We follow the same convention and build an anomaly dataset out of the NGSIM dataset [11, 12].
For assessment, we use the conventional metrics namely the Area Under The Receiver Operating Characteristic (AUC-ROC), the Area Under The Precision Recall Curve (AUC-PR), and the Equal Error Rate.
VI-A Adversarial Anomaly Trajectory Generation
Building upon discussions from prior sections, the absence of dedicated highway-based trajectory anomaly datasets presents a challenge in validating our anomaly detection methodologies. To circumvent this limitation, we adopt the recent advancements in adversarial anomaly generation as a benchmark for our proposed approach. Zhang et al., in their work [68], introduce an adversarial attack-based technique designed to craft realistic anomalous trajectories by perturbing standard trajectories within a dataset.
[68] uses six metrics to generate anomalies, such as ADE, FDE, deviation towards the left and right side of the lateral direction, and deviation towards the front and rear side of the longitudinal direction. The metrics for optimizing targeted disruptions are given as [24]:
$D(\alpha,R)=(p_{\alpha}-s_{\alpha})^{T}\cdot R(s_{\alpha+1},s_{\alpha}),$ (11)
where $\alpha$ represents the time frame, $R$ is a function that produces the unit vector towards a designated direction (lateral direction in our case), and $p$ and $s$ symbolize the predicted and actual vehicle positions, respectively.
Based on metrics, optimization has six different objective functions with several hard constraints based on the characteristics of the road and vehicle movements discussed in [68]. Finally, the optimization problem is designed to satisfy the objective and obey the constraints. To ensure adherence to the constraints, we adjust the perturbation, denoted ( $\Delta$), based on the historical trajectory ($H_{i}$) of the target vehicle $i^{th}$ and the constraint function ($C$). The optimization problem, as shown in the equation below, determines the maximum coefficient $\theta_{max}$ that scales the perturbation while satisfying all constraints for the vehicle of interest.
To evaluate our anomaly detection methodology, we produce anomalous data for the NGSIM dataset [12, 11] using the described adversarial trajectory generation technique.
VI-B Anomaly Detection Results
As discussed, the output of VT-Former is compared with the actual trajectory to classify each trajectory as normal or anomalous. As shown in Fig. 3, the results demonstrate an AUC-ROC of 0.71 and AUC-PR of 0.82 with an EER of 0.32. The results of the driving anomaly detector demonstrate moderately effective performance, with an AUC-ROC score of 0.71, indicating a reasonable ability to distinguish anomalies from non-anomalies. In contrast, the AUC-PR score of 0.82 reflects a strong balance between precision and recall in identifying anomalies. The EER at 0.32 signifies a specific threshold point where the detector equally balances false positives and false negatives. These findings suggest that the model is proficient in identifying anomalies and minimizing false positives, making it well-suited for applications where precision is crucial, such as driving anomaly detection in security or maintenance contexts. Fine-tuning the operating threshold may further optimize the model’s performance to align with specific application requirements.
VII Conclusion
This work introduced VT-Former, an innovative trajectory prediction algorithm that leverages transformer architectures in conjunction with a novel tokenization scheme utilizing graph attention mechanisms denoted as Graph Attentive Tokenization (GAT). Extensive experimentation was conducted across three benchmark datasets illustrating SotA performance results, as well as assessing the model’s generalizability and robustness. Further experimentation aimed to demonstrate the real-time capabilities of VT-Former when deployed on an embedded platform, in addition to evaluating its performance under shorter observed horizons. Moreover, the study extended its contributions by implementing a driving anomaly detection approach built upon the foundation of VT-Former, thereby exemplifying a practical application of the proposed model.
Acknowledgment
This research is supported by the National Science Foundation (NSF) under Award No. 1932524.